The power rating of a water pump is 4200 watts, if it is run for 600s, how many joules of work will it do in that time?


1
Expert's answer
2022-02-10T08:36:46-0500

Given:

"P=4200\:\rm W"

"t=600\:\rm s"


The work done by a power pump

"W=P*t\\
=\rm 4200\: W*600\: s=2520000\: J=2.52\: MJ"